# Scene-Adaptive Motion Planning with Explicit Mixture of Experts and Interaction-Oriented Optimization

### v2

@@ Abstract
Despite over a decade of development, autonomous driving trajectory planning in complex urban environments continues to encounter significant challenges. These challenges include the difficulty in accommodating the multi-modal nature of trajectories, the limitations of single expert model in managing diverse scenarios, and insufficient consideration of environmental interactions. To address these issues, this paper introduces the EMoE-Planner, which incorporates three innovative approaches. Firstly, the Explicit MoE (Mixture of Experts) dynamically selects specialized experts based on scenario-specific information through a shared scene router. Secondly, the planner utilizes scene-specific queries to provide multi-modal priors, directing the model’s focus towards relevant target areas. Lastly, it enhances the prediction model and loss calculation by considering the interactions between the ego vehicle and other agents, thereby significantly boosting planning performance. Comparative experiments were conducted on the Nuplan dataset against the state-of-the-art methods. The simulation results demonstrate that our model consistently outperforms SOTA models across nearly all test scenarios. Our model is the first pure learning model to achieve performance surpassing rule-based algorithms in almost all Nuplan closed-loop simulations.
Index Terms:
Autonomous Driving, Trajectory Planning, Imitation Learning, Mixture of Experts.
@@ I Introduction
Conversely, data-driven approaches, including imitation learning [4, 5, 6, 7, 8], address edge cases like extreme weather and complex traffic, capturing human-like driving behaviors from expert data. Reinforcement learning  [9, 10, 11] enables dynamic optimization through advanced reward mechanisms. These systems offer lower costs and faster iterations compared to rule-based alternatives.
@@
  • Propose an explicit Mixture of Experts (EMoE) architecture that employs a shared scene router for explicit scenario classification and selects corresponding experts based on the classification results. This approach simplifies the responsibilities and learning complexity of each expert, thereby reducing training difficulty and enhancing model performance.
  • Propose a novel scene-specific learnable query to enable the model to focus on area that the goal point is located. The query simultaneously provides prior information to the model at both the decision level and the modality level, allowing the model to achieve better training efficiency and performance with only a small number of queries.
  • Introduce an Interaction-Oriented Loss that explicitly accounts for the interaction between the ego vehicle and surrounding agent within the loss function. Furthermore, the prediction of surrounding agents trajectories considers their interaction with the ego vehicle, which not only improves the prediction accuracy but also indirectly optimizes the scene encoding.
@@ II Related Work
Multi-modality is a significant and persistent challenge for autonomous driving trajectory planning. Early approaches [14, 15, 16] overlooked the inherent multi-modal nature of driving scenarios, particularly anchor-free methods that relied solely on environmental and map information to predict future trajectories. Even recent end-to-end models like UniAD [12], specifically its motion former component, leverage map and object information to generate trajectories via anchor-free motion queries. However, these approaches risk modal collapse due to their limited capacity to represent diverse potential behaviors. Current state-of-the-art models are increasingly adopting anchor-based strategies. PLUTO [4] and MPNP [17] employ a semi-anchor approach, leveraging reference lines as lateral anchors and anchor-free queries longitudinally to approximate multi-modality. While achieving impressive performance, these methods are heavily dependent on accurate road feature and topology information. Critically, existing perception models, including end-to-end architectures, struggle to directly output reliable reference line information, requiring the addition of rule-based layers for generation. This reliance on rule-based preprocessing hinders practical deployment and end-to-end learning. Alternatively, VAD2 [18] and Hydra MDP [19] utilize pre-generated trajectory libraries as anchors to cover a wide range of possible trajectory modalities. These methods select the optimal trajectory by scoring candidates. Further refinements, such as applying Gaussian Mixture Models (GMMs) to trajectory generation [20, 21] or sampling trajectories from predicted heatmaps [22, 23], aim to improve the distribution rationality and selection probability of generated trajectories. Multi-modal Motion[24] also leverages generated trajectories as anchors but introduces a region-specific training paradigm, selecting different regional anchors based on ground truth. Diffusion models are gaining traction in trajectory generation [25, 26] due to their inherent multi-modal capabilities. However, their accuracy, efficiency, and stability require further enhancement. Approaches that leverage generated trajectories as anchors involve a large number of anchor points and impose excessive constraints on the diversity of the final trajectories.
@@
Consequently, researchers have explored using trajectory endpoints or destination points as anchors [27, 28, 29, 30]. MTR [29] directly generates trajectory endpoints by clustering ground truth data, leveraging endpoints as intent points with learnable position embeddings serving as static intent queries. GoalFlow [30], builds upon these approaches by directly generating trajectory endpoints as anchors. While endpoint anchors enhance trajectory interpretability and preserve multi-modality, current methods use a single anchor set for all scenes. Dense anchor sets increase computation, while sparse anchor sets limit modality. This paper introduces a model with scene-specific endpoint anchors, balancing anchor density and modality representation across diverse driving environments.
@@
Furthermore, the complexity of urban scenarios presents unique challenges, as differences between scenes persist despite some common planning rules. STR2 [31] address this by replacing the GPT2 backbone in STR [32] with an MoE architecture, substituting the feedforward network in the Transformer [33] with an MoE layer. This adaptation leverages MoE’s classification capabilities, significantly boosting performance in closed-loop simulations. MoE [34, 35], widely used in large language models like deepseek-v2 [36], Qwen1.5 [37], Switch Transformer [38] and Skywork-MoE [39], is apt for planning tasks that require scenario-specific strategies. However, traditional MoE’s unsupervised expert selection often leads to imbalanced loads and reduced interpretability. While techniques like auxiliary load balancing losses, as proposed in Switch Transformer [38] and DeepSeek-V2 [36], can mitigate this issue, they also distort the overall loss landscape and potentially compromise model generalization. Therefore, this paper introduces an improved MoE architecture tailored for trajectory planning, which is not only highly interpretable but also provides valuable insights into data collection strategies based on the performance of individual experts.
@@
Modeling multi-agent interactions remains a critical challenge in autonomous driving trajectory planning. The mutual influence between the planned trajectory and predicted trajectories presents fundamental complexity. Prediction-first paradigms [40, 12, 41] sequentially predict agent trajectories then plan ego motion, but neglect the ego’s impact on others, often yielding over-conservative behaviors. Planning-first paradigms [42, 43] conversely plan ego trajectories before predicting agent responses, risking over-aggressive maneuvers through unrealistic collision-avoidance assumptions. Joint optimization frameworks [44, 45] simultaneously optimize both trajectories with shared objectives, yet suffer from prediction-execution mismatches when agent cooperation assumptions violate real-world responses. Iterative refinement [46, 47], while theoretically optimal for capturing interaction dynamics, incurs prohibitive computational costs for real-time applications. To fully consider the influence of interaction, this paper adapts planning-first paradigm and introduces an interactive regression loss that penalizes trajectory deviations caused by mutual ego-agent influences during optimization.
@@ III-A Overview
Figure 1 illustrates the framework of the proposed model. The model receives the ego vehicle state information $F_{e}\in R^{D_{e}}$, dynamic object information $F_{a}\in R^{N_{a}\times T_{h}\times D_{a}}$, static object information $F_{s}\in R^{N_{s}\times D_{s}}$, and map information $F_{m}\in R^{N_{m}\times L_{p}\times D_{m}}$. These inputs are processed through respective encoding modules to obtain D-dimensional features. The details of these input encoding modules will be elaborated in Section 3.2. It is worth noting that the positional and heading angle information in the input data are relative values defined in the ego vehicle’s coordinate system.
@@ III-C Explicit MOE
In our decoder, we incorporate a MoE model, assigning specialized experts to address distinct driving scenarios. In autonomous driving trajectory planning tasks, different scenarios often contain varying types of information, requiring the ego vehicle to focus on specific aspects depending on the situation. Tailoring specific rules for each type of scenarios simplifies them considerably compared to universal rules. By employing specialized experts for varying scenarios in our planning model, we streamline the learning demands for each expert, enhancing the efficiency of the learning process.
Our approach deviates from vanilla MoE models by introducing an architecture designed specifically for our task. Figure 2(a) depicts a standard MoE model where at each layer, a router selects top $N_{E}$ best-suited expert from all experts based on the input information, with $N_{E}=1$ in our case, to generate the output. However, this architecture has several limitations: the router lacks explicit constraints, making its selection logic difficult to interpret and prone to confusion in similar scenarios. It also risks an uneven distribution of expert utilization.
Figure 2(b) introduces the explicit MoE (EMoE) architecture developed in this study. Our architecture distinguishes itself by featuring an independent scene router, separate from the expert networks, with a consistent router utilized across all layers. The router inputs are scene features rather than traditional MoE inputs. In trajectory planning tasks, the task scenarios can be explicitly divided into seven categories: turn left at intersections, go straight at intersections, turn right at intersections, go straight at non-intersections, roundabouts, U-turns, and others. Training data is tagged with scene labels, which supervise the output of the scene router. The unified scene router in our EMoE model ensures that the same expert index is selected for each layer when the scenario type remains consistent. This alignment accelerates model convergence. In contrast, traditional MoE models suffer from reduced training efficiency as they require continuous expert selection due to varying inputs at each layer.
@@
A further benefit of EMoE is the precise control over input data distribution, allowing for a more focused approach to data collection. Notably, less frequent scenarios such as U-turns and roundabouts may yield poorer expert performance due to limited data. To address this, targeted data augmentation for these specific scenarios can be implemented to enhance expert effectiveness.
@@ III-D Scene-specific query
We introduce scene-specific queries that incorporate prior knowledge at both decision and modality levels, enhancing the model’s focus on future areas relevant to intended actions. This approach not only stabilizes training but also simplifies the learning process. This concept aligns with human driving behaviors where attention is naturally directed towards areas pertinent to intended maneuvers, such as the left-front region when preparing to turn left. As discussed in section 3.3, driving scenarios are categorized into seven types. Utilizing a well-organized dataset, we apply the k-means clustering algorithm to cluster the endpoints of ground-truth trajectories for each scenario type. This results in a set of scene-specific target points denoted as $G\in R^{7\times K_{a}\times 2}$, with $K_{a}$ representing the number of clusters and the second dimension representing x and y coordinates. These points, $G$, serve as anchors analogous to endpoint anchors in existing methodologies. Following scenario classification by the scene router, a specific set of target points $G_{i}\in R^{K_{a}\times 2}$is selected for each scene, facilitating more precise model training and prediction accuracy.
@@
Where $Q_{\text{learnable}}\in R^{K_{a}\times D}$ is a learnable query. Each $Q_{\text{mode}}$ integrates prior knowledge across scenario and modality levels, targeting specific trajectory predictions within designated regions for each scenario. This architecture promotes trajectory generation with multiple modalities while ensuring concentrated attention, thereby stabilizing training and enhancing convergence speed. Moreover, the efficiency of our focused attention mechanism allows for a reduced number of queries (24 in our experiments), in contrast to traditional anchor-based models, leading to more effective region coverage and improved prediction capabilities. Figure 3 shows the distribution of anchor points across the seven scenarios used in our study.
@@ III-E1 Interaction-oriented loss
In existing imitation learning frameworks [4, 5, 6], trajectory regression loss is typically calculated assuming equal importance across all trajectory points, with uniform weighting. However, real-world driving behaviors demonstrate a prioritization of immediate, short-term actions. Nearby obstacles, for instance, may necessitate evasive maneuvers and significantly influence the trajectory, whereas distant obstacles might not alter driving behavior significantly unless they come into closer proximity. This suggests the need for a differentiated approach to learning expert trajectories, where immediate responses to the near-future environment are prioritized over responses to distant conditions.
@@
For the i-th agent, the process of $t_{\text{in}}^{i}$ and $t_{\text{out}}^{i}$ is shown in Equation 8-9. The interaction interval is defined as the union of all ($t_{\text{in}}^{i}$,$t_{\text{out}}^{i}$) intervals, as shown in Equation 10.
@@
$(t_{\text{in}},t_{\text{out}})=\bigcup_{i}(t_{\text{in}}^{i},t_{\text{out}}^{i})$ (10)
@@
$w=\begin{cases}1&t_{\text{in}}\leq t\leq t_{\text{out}}\\ e^{-k_{R}t}&\text{else}\end{cases}$ (11)
@@
where $w_{1}$,$w_{2}$,$w_{3}$, $w_{4}$ and $w_{5}$ are weights of each loss. $\mathcal{L}_{cls}$, $\mathcal{L}_{col}$ and $\mathcal{L}_{pred}$ are imitation class loss, collision loss and prediction loss. These losses are similar to those in existing work [4, 5], thus we will not introduce them in detail. $\mathcal{L}_{router}$ is the loss of scene router mentioned in section 3.3. It can be calculated with the following equation.

### v3

@@ Abstract
Despite over a decade of development, autonomous driving trajectory planning in complex urban environments continues to encounter significant challenges. These challenges include the difficulty in accommodating the multi-modal nature of trajectories, the limitations of the single expert model in managing diverse scenarios, and insufficient consideration of environmental interactions. To address these issues, this paper introduces the EMoE-Planner, which incorporates three innovative approaches. Firstly, the Explicit MoE (Mixture of Experts) dynamically selects specialized experts based on scenario-specific information through a shared scene router. Secondly, the planner utilizes scene-specific queries to provide multi-modal priors, directing the model’s focus towards relevant target areas. Lastly, it enhances the prediction model and loss calculation by considering the interactions between the ego vehicle and other agents, thereby significantly boosting planning performance. Comparative experiments were conducted on the Nuplan dataset against the state-of-the-art methods. The simulation results demonstrate that our model consistently outperforms SOTA models across nearly all test scenarios. Our model is the first pure learning model to achieve performance surpassing rule-based algorithms in almost all Nuplan closed-loop simulations.
@@ I Introduction
Conversely, data-driven approaches, including imitation learning [6, 7, 20], address edge cases like extreme weather and complex traffic, capturing human-like driving behaviors from expert data. Reinforcement learning [19, 21] enables dynamic optimization through advanced reward mechanisms. These systems offer lower costs and faster iterations compared to rule-based alternatives.
@@
  • Propose an Explicit Mixture of Experts (EMoE) that replaces the black-box routing in traditional MoE with a semantically meaningful scene classifier. This explicit routing mechanism assigns well-defined tasks to each expert, thereby reducing their individual learning complexity. Moreover, it ensures inherent load balance by allowing direct control over the training data distribution across different scenario types.
  • Introduce a scene-specific query that dynamically adapts to the scenario type identified by EMoE architecture. The scene-specific query confines the output space to only relevant maneuvers (e.g., left turns for a left-turn scenario), thereby raising the quality floor by avoiding the generation of unrealistic or poor-quality trajectories.
  • Introduce an Interaction-Oriented Loss to steer the model’s attention toward what is more important within an interaction. Furthermore, the prediction of surrounding agents trajectories considers their interaction with the ego vehicle, which not only improves the prediction accuracy but also indirectly optimizes the scene encoding.
@@ II Related Work
Multi-modality is a significant and persistent challenge for autonomous driving trajectory planning. Early approaches [43, 8] overlooked the inherent multi-modal nature of driving scenarios, particularly anchor-free methods that relied solely on environmental and map information to predict future trajectories. Even recent end-to-end models like UniAD [17], specifically its motion former component, leverage map and object information to generate trajectories via anchor-free motion queries. However, these approaches risk modal collapse due to their limited capacity to represent diverse potential behaviors. Current state-of-the-art models are increasingly adopting anchor-based strategies. PLUTO [6] employ a semi-anchor approach, leveraging reference lines as lateral anchors and anchor-free queries longitudinally to approximate multi-modality. Despite their impressive performance, these methods are limited by a dependence on detailed maps. Since perception models cannot yet reliably output reference lines directly, rule-based layers are still required, which complicates deployment and prevents full end-to-end learning. Alternatively, VAD2 [5] and Hydra MDP [27] utilize pre-generated trajectory libraries as anchors to cover a wide range of possible trajectory modalities. These methods select the optimal trajectory by scoring candidates. Further refinements, such as applying Gaussian Mixture Models (GMMs) to trajectory generation [4, 16] or sampling trajectories from predicted heatmaps [14, 15], aim to improve the distribution rationality and selection probability of generated trajectories. Diffusion models are gaining traction in trajectory generation [24, 28, 48] due to their inherent multi-modal capabilities. By dividing the action space with predefined anchor points and employing a truncated diffusion policy, DiffusionDrive [28] generates multimodal trajectories ten times faster than traditional diffusion strategies. However, the generated trajectories lack adequate constraints, making it challenging to balance high quality with diversity, and in some cases, it may even produce trajectories of very poor quality. To address this, DiffusionPlanner [48] augments a pre-trained diffusion model with classifier guidance, which applies constraints to the trajectories. This ensures both the quality and diversity of the generated paths while effectively avoiding low-quality results.
@@
Consequently, researchers have explored using trajectory endpoints or destination points as anchors [46, 35, 42]. MTR [35] directly generates trajectory endpoints by clustering ground truth data, leveraging endpoints as intent points with learnable position embeddings serving as static intent queries. GoalFlow [42], builds upon these approaches by directly generating trajectory endpoints as anchors. While endpoint anchors enhance trajectory interpretability and preserve multi-modality, current methods use a single anchor set for all scenes. Dense anchor sets increase computation, while sparse anchor sets limit modality. This paper introduces a model with scene-specific endpoint anchors, balancing anchor density and modality representation across diverse driving environments.
@@
Furthermore, the complexity of urban scenarios presents unique challenges, as differences between scenes persist despite some common planning rules. STR2 [37] address this by replacing the GPT2 backbone in STR [38] with an MoE architecture, substituting the feedforward network in the Transformer [41] with an MoE layer. This adaptation leverages MoE’s classification capabilities, significantly boosting performance in closed-loop simulations. MoE [22, 26], widely used in large language models (LLM) like deepseek-v2 [29], Qwen1.5 [2] and Switch Transformer [12], is apt for planning tasks that require scenario-specific strategies. However, traditional MoE’s unsupervised expert selection often leads to imbalanced loads and reduced interpretability. While techniques like auxiliary load balancing losses, as proposed in Switch Transformer [12] and DeepSeek-V2 [29], can mitigate this issue, they also distort the overall loss landscape and potentially compromise model generalization. Therefore, this paper introduces an improved MoE architecture tailored for trajectory planning, which is not only highly interpretable but also provides valuable insights into data collection strategies based on the performance of individual experts.
@@
Modeling multi-agent interactions remains a critical challenge in autonomous driving trajectory planning. The mutual influence between the planned trajectory and predicted trajectories presents fundamental complexity. Prediction-first paradigms [44, 17, 23] sequentially predict agent trajectories then plan ego motion, but neglect the ego’s impact on others, often yielding over-conservative behaviors. Planning-first paradigms [1, 34] conversely plan ego trajectories before predicting agent responses, risking over-aggressive maneuvers through unrealistic collision-avoidance assumptions. Joint optimization frameworks [47, 31] simultaneously optimize both trajectories with shared objectives, yet suffer from prediction-execution mismatches when agent cooperation assumptions violate real-world responses. Iterative refinement [32, 13], while theoretically optimal for capturing interaction dynamics, incurs prohibitive computational costs for real-time applications. GameFormer [20] employs a layered architecture in which each layer refines its result based on predictions from the previous layer, thereby balancing computational cost and performance by adjusting the number of layers. To fully consider the influence of interaction, this paper adapts planning-first paradigm and introduces an interactive regression loss that penalizes trajectory deviations caused by mutual ego-agent influences during optimization.
@@ III-A Overview
Figure 1 illustrates the framework of the proposed model. The model receives the ego vehicle state information $F_{e}\in R^{D_{e}}$, dynamic object information $F_{a}\in R^{N_{a}\times T_{h}\times D_{a}}$, static object information $F_{s}\in R^{N_{s}\times D_{s}}$, and map information $F_{m}\in R^{N_{m}\times L_{p}\times D_{m}}$. These inputs are processed through respective encoding modules to obtain D-dimensional features. The details of these input encoding modules will be elaborated in Section III-B. It is worth noting that the positional and heading angle information in the input data are relative values defined in the ego vehicle’s coordinate system.
@@ III-C Explicit MoE
In autonomous driving trajectory planning, different scenarios require the ego vehicle to focus on specific infomation. In rule-based planners, tailoring specific rules for each type of scenarios simplifies them considerably compared to universal rules. Inspired by this, employing specialized experts for varying scenarios in a planning model can similarly simplify each expert’s learning demands and enhance efficiency. To this end, we integrate a MoE structure into our model design.
Different from vanilla MoE, we propose Explicit MoE (EMoE), a new variant tailored for our task. Figure 2(a) depicts a standard MoE model where at each layer of the decoder, a router selects top $N_{E}$ expert and their outputs are combined through weighted summation to produce the final output. However, this architecture has several limitations. Firstly, the router operates as a black box, whose outputs lack semantic meaning, making the selection logic neither interpretable nor monitorable. Secondly, it is challenging to ensure proper load balance among the different experts without extra design or loss. In LLMs, tasks are often ambiguous and lack clear delineation, making a semantic-agnostic router appropriate. This allows the model to learn implicit task classification and expert selection. However, for autonomous driving planning, scenarios can be explicitly divided into a finite set of sub-tasks. Thus, binding each expert to a specific sub-task emerges as a natural choice.
Figure 2(b) illustrates EMoE. The architecture is distinguished by using only one router across all decoder layers. This router takes scene features as input, and its output, namely the identified scenario type, is shared by every layer. In this paper, the task scenarios can be explicitly divided into seven categories: turn left in junctions, go straight in junctions, turn right in junctions, go straight, roundabouts, U-turns, and others. The ”others” category encompasses complex scenarios such as lane merging/diverging, lane number changes, and pull-over/starting, etc., which are difficult to classify reliably under conventional rules. All training data are pre-labeled using fixed rules. These data labels are then used to supervise the training of router. This enables the router to learn explicit and fixed routing rules, while each expert is dedicated to processing one specific scenario (e.g., Expert 3 only handles right-turns). Consequently, each expert specializes more effectively, improving individual performance—especially for data-scarce scenarios like U-turns and roundabouts—and boosting overall model capability. Notably, experts sharing the same index across all layers are assigned to the same scenario (e.g., Expert 3 in all layers process right-turn scenarios), unlike in vanilla MoE where experts in different layers are independent. This design also accelerates model convergence.
@@
Furthermore, EMoE inherently enables load balancing via its data-centric design: each expert’s workload is determined solely by the proportion of its assigned scenario in the dataset. This allows for targeted fine-tuning of experts by strategically adjusting the data distribution during training.
@@ III-D Scene-specific query
We introduce scene-specific queries that incorporate prior knowledge at both decision and modality levels, enhancing the model’s focus on future areas relevant to intended actions. This approach not only stabilizes training but also simplifies the learning process. This concept aligns with human driving behaviors where attention is naturally directed towards areas pertinent to intended maneuvers, such as the left-front region when preparing to turn left. As discussed in section III-C, driving scenarios are categorized into seven types based on a set of predefined rules. Every training sample is then annotated with one of these scenario labels. By extracting all ground-truth trajectories from a given scenario type, we obtain a set that nearly covers all possible motion modalities of the ego vehicle over the 8-second horizon within that scenario. The endpoints of these trajectories represent the possible vehicle positions at 8 seconds. We then cluster all endpoints into $K_{a}$ classes using k-means, and the resulting cluster centroids serve as the anchor points for that scenario. Repeating this process for all seven scenario types yields a set of scene-specific anchor points, denoted as $G\in R^{7\times K_{a}\times 2}$ where the last two dimensions correspond to the x and y coordinates, respectively. Following scenario classification by the scene router, a specific set of target points $G_{i}\in R^{K_{a}\times 2}$ is selected for each scene, facilitating more precise model training and prediction accuracy.
@@
Where $Q_{\text{learnable}}\in R^{K_{a}\times D}$ is a learnable query. Each $Q_{\text{mode}}$ integrates prior knowledge across scenario and modality levels, targeting specific trajectory predictions within designated regions for each scenario. This architecture promotes trajectory generation with multiple modalities while ensuring concentrated attention, thereby stabilizing training and enhancing convergence speed. Moreover, the efficiency of our focused attention mechanism allows for a reduced number of queries (24 in our experiments), in contrast to traditional anchor-based models, leading to more effective region coverage and improved prediction quality. Figure 3 shows the distribution of anchor points across the seven scenarios used in our study.
@@ III-E1 Interaction-oriented loss
In existing imitation learning frameworks [6, 7], trajectory regression loss is typically calculated assuming equal importance across all trajectory points, with uniform weighting. However, real-world driving behaviors demonstrate a prioritization of immediate, short-term actions. Nearby obstacles, for instance, may necessitate evasive maneuvers and significantly influence the trajectory, whereas distant obstacles might not alter driving behavior significantly unless they come into closer proximity. This suggests the need for a differentiated approach to learning expert trajectories, where immediate responses to the near-future environment are prioritized over responses to distant conditions.
@@
For the i-th agent, the process of $t_{\text{in}}^{i}$ and $t_{\text{out}}^{i}$ is shown in Equation 8-9. The interaction interval is defined as the union of all ($t_{\text{in}}^{i}$$t_{\text{out}}^{i}$) intervals, as shown in Equation 10.
@@
$(t_{\text{in}},t_{\text{out}})=\bigcup_{i}(t_{\text{in}}^{i},t_{\text{out}}^{i})\vskip-5.69046pt$ (10)
@@
$W_{t}=\begin{cases}1&t_{\text{in}}\leq t\leq t_{\text{out}}\\ e^{-k_{R}t}&\text{else}\end{cases}$ (11)
@@
where $w_{1}$$w_{2}$$w_{3}$, $w_{4}$ and $w_{5}$ are weights of each loss. $\mathcal{L}_{cls}$, $\mathcal{L}_{col}$ and $\mathcal{L}_{pred}$ are imitation class loss, collision loss and prediction loss. These losses are similar to those in existing work [6, 7], thus we will not introduce them in detail. $\mathcal{L}_{router}$ is the loss of scene router mentioned in section 3.3. It can be calculated with the following equation.